OpenStudy (jessiegonzales):

Solve 1/3x-5 - 5/x-2= 12

OpenStudy (anonymous):

Is the equation:\[\frac{1}{3x-5} - \frac{5}{x-2} = 12\]

OpenStudy (jessiegonzales):

yes :)

OpenStudy (anonymous):

Start by subtracting the two fractions as you would do a normal fraction, by cross multiplying.

OpenStudy (jessiegonzales):

(x-2)-(15x-30)

OpenStudy (anonymous):

Great! That is the numerator and the denominator would be?

OpenStudy (jessiegonzales):

x=29/18 ??

OpenStudy (anonymous):

1/(3x-5) - 5/(x-2) - 12//(3x-5)(x-2) = 0

OpenStudy (jessiegonzales):

A. x=34/9 B. x=-29/18 C. x=-34/9 D. x=29/18

OpenStudy (anonymous):

((X-2)-5(3x-5)-12)/(3x-5)(x-2)=0

OpenStudy (anonymous):

It goes like this: \[\frac{(x-2)-(15x-25)}{(3x-5)(x-2)} = 12\] \[\frac{x-2-15x+25}{(3x-5)(x-2)} = 12\] \[\frac{-14x+23}{(3x-5)(x-2)} = 12\] \[-14x+23 = 12(3x-5)(x-2)\] Can you solve further?

OpenStudy (jessiegonzales):

Is the answer D?

OpenStudy (anonymous):

Sorry 12 instead of 0 ((X-2)-5(3x-5)-12)/(3x-5)(x-2)=12

OpenStudy (anonymous):

The answer does not match any of the four given choices. Check the question again?

OpenStudy (jessiegonzales):

@Chad123 pls is it D???

OpenStudy (anonymous):

@jessiegonzales if all you care about is the answer, then you should learn the easiest way

OpenStudy (anonymous):

Plug in each of your answer choices until one of them is correct

OpenStudy (anonymous):

And that works for every question...